Potassium hydrogen phthalate is a solid, monoprotic acid frequently used in the laboratory to standardize strong base solutions. It has the unwieldy formula of KHC8H4O4. This is often written in shorthand notation as KHP. What volume of a 0.299 M calcium hydroxide solution is needed to exactly neutralize 3.24 grams of KHP ? ANSWER: mL calcium hydroxide
Answer #1

The reaction followed is,

2 KHP + Ca(OH)2   2 H2O + Ca(KP)2

Molar mass of KHP = 204.22 g/mol

number of moles of KHP in 3.24 grams of KHP =given mass / molar mass = (3.24 / 204.22) = 0.0159 mol

From the reaction we see that 2 moles of KHP reacts with 1 mol of Ca(OH)2

or, 1 mol of KHP reacts with 0.5 mol of Ca(OH)2

mathematically,

1 mol of KHP = 0.5 mol of Ca(OH)2  

Thus, by unitary method

0.0159 mol of KHP = 0.5 * 0.0159 = 7.95 * 10-3 mol of Ca(OH)2  

We are provided with Ca(OH)2 solution with concentration 0.299 M, or in other words

0.299 mol of Ca(OH)2 in 1L solution

mathematically,

0.299 mol of Ca(OH)2 = 1L solution

0.299 mol of Ca(OH)2 = 1000 mL solution

By unitary method,

1 mol of Ca(OH)2 = (1000 / 0.299) mL solution

Thus the required amount of calcium hydroxide solution is,

7.95 * 10-3 mol of Ca(OH)2 = 7.95 * 10-3 * (1000 / 0.299) mL solution = 26.59 mL

Therefore to neutralize 3.24 g of KHP , 26.59 mL of 0.299 M Ca(OH)2solution is required.